# Adv-BMT: Bidirectional Motion Transformer for Safety-Critical Traffic Scenario Generation

### v1

@@ A.1 BMT Architecture
Scenario Format and Preprocessing
The model takes real-world scenarios with the format from ScenarioNet li2023scenarionetopensourceplatformlargescale , which unifies real-world datasets including caesar2020nuscenesmultimodaldatasetautonomous  caesar2022nuplanclosedloopmlbasedplanning  wilson2023argoverse2generationdatasets .
@@
The model performs a series of preprocessing on both static map features and dynamic agent trajectories. We compute the global boundary to extract a consistent map center and heading. Each map feature is decomposed into a sequence of vectorized segments, where each vector is represented by a start and end point in 3D coordinates. These vectors are then augmented with directional relative positions, segment headings, and lengths. Semantic attributes are binary indicators that encode feature types. The map feature types (lane, crosswalk, broken line, yellow line, stop sign, etc.) are encoded for semantic information. Features are then centralized to map center. Traffic light states are also encoded and aligned with the steps. Similarly, agent trajectories are also centralized for agent feature encode, with information of (position, heading, velocity, shape, and type) extracted across time for temporal sequences. Besides, there is reordering for ego agent and agents to be predicted. We extract a 16-dimensional state vector at each timestep. The preprocessing ensures all trajectories are represented in a consistent spatial frame and translated to an ego-centric coordinate system.
Tokenization
@@
We formulate motion prediction as an auto-regressive next-token prediction problem, where each motion token corresponds to a discretized control input for a fixed time interval. Tokenization process maps continuous agent motion (acceleration and yaw rate) into discrete 2D bins. The tokenization process considers candidate tokens sampled from a fixed grid of bins, and simulates the resulting motion over a short duration, and then selects the best-matching action by minimizing the contour alignment error between the predicted agent shape and the ground-truth position and heading. We adopt a simplified version of the bicycle model to parameterize agent motion using longitudinal acceleration and yaw rate within predefined bounds: acceleration is limited to the range of $[-10,10]$ m/s2, and the yaw rate is constrained to $[-\pi/2,\pi/2]$ rad/s. With predicted motion token sequences, the trajectory can be reconstructed by mapping the token back to the acceleration and yaw angle change.
@@
In both forward and reverse directions, motion tokens are decoded into continuous trajectories using the same tokenizer. Given an initial state $\tau_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, the forward decoding process simulates the agent’s next state $\tau_{t+1}$ by applying a tokenized control action $z_{t}=(a_{t},\delta_{t})\in\mathcal{A}$, where $\mathcal{A}$ denotes the discrete token space. This is repeated autoregressively over the sequence of predicted tokens to reconstruct the full trajectory $\boldsymbol{\tau}_{t:t+T}$. In contrast, reverse decoding starts from a known future state $\tau_{t+1}$, the model evaluates all possible token candidates $z_{t}\in\mathcal{A}$, simulates their inverse dynamics using $\Delta t\to-\Delta t$, and selects the token that best reconstructs the preceding state $\tau_{t}$. The ability to operate in both directions is a key distinction of our approach: forward prediction enables open-loop simulation of future behaviors, while reverse prediction allows us to trace back from a desired outcome (e.g., a collision state) to plausible initiating actions.
Decoding Architecture.
@@
The decoder follows a GPT-like structure composed of stacked cross-attention layers. Each layer integrates three structured attention modules: agent-to-agent (A2A), agent-to-temporal (A2T), and agent-to-scene (A2S). These modules attend over dynamically constructed graphs defined by spatial or temporal adjacency. Relational information across modalities is captured via multiple embeddings: Agent-to-Agent Relation Embedding, Agent-to-Time Relation Embedding, and Agent-to-Scene Relation Embedding, each encoding context-specific spatial information. For token construction, several embeddings are used, including the Agent Type Embedding, Agent Shape Embedding, Agent ID Embedding, and the Motion Token Embedding which maps discrete control tokens. A Continuous Motion Feature Embedding is applied to embed acceleration and yaw rate attributes. Auxiliary embeddings include the Special Token Embedding for indicating sequence boundaries, and the Backward Prediction Indicator Embedding to distinguish between forward and backward prediction modes. For each attention edge, a relative relation embedding is computed using a Fourier encoder and added to the key and/or value vectors. The input agent token $\mathbf{X}\in\mathbb{R}^{B\times T\times N\times d}$ (batch size $B$, time steps $T$, number of agents $N$, hidden dimension $d$) is progressively updated across layers by aggregating contextual features from other agents, their temporal history, and relevant map elements.
Motion Token Embedding
@@
Prediction Heads and Outputs
@@
where $\phi(\cdot)$ denotes the GELU activation and $K^{2}$ is the number of discrete motion tokens (from a $K\times K$ acceleration–yaw bin grid).
The resulting logits are used to predict the next motion token at each time step. During inference, we generate motion tokens using nucleus (top-$p$) sampling.
@@
Our model makes a prediction in the interval of 5 time steps (0.5 seconds). To simulate the effect of a motion token over a fixed time step $\Delta t=0.5\,\mathrm{s}$, we adopt midpoint integration based on a simplified bicycle model. In forward predictions, given a current state $\mathbf{s}_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, the model computes the next speed and heading as $v_{t+1}=v_{t}+a\cdot\Delta t$ and $\theta_{t+1}=\theta_{t}+\omega\cdot\Delta t$. The average speed and heading are then defined as $\bar{v}=\frac{v_{t}+v_{t+1}}{2}$ and $\bar{\theta}=\left(\frac{\theta_{t}+\theta_{t+1}}{2}\right)$. In reverse prediction, the process is reverted. In the backward direction, the process is inverted. Given a future state $\mathbf{s}_{t+1}$, the model enumerates all possible token candidates and inverts the dynamics: $v_{t}=v_{t+1}-a\cdot\Delta t$ and $\theta_{t}=\theta_{t+1}-\omega\cdot\Delta t$. The average quantities $\bar{v}$ and $\bar{\theta}$ are computed similarly and used to derive the previous position:
@@
Decoder Training Loss
@@ A.2 Training Details
Appendix B Experiment Details
Training Environment
@@ Appendix C Quantitative Results
C.1 Waymo Open Sim Agents Challenge
@@
We evaluated BMT resutls on 400 WOMD validation scenarios using Waymo Open Sim Agents Challenge (WOSAC) 2024 montali2023waymoopensimagents . Evaluation results are summarized in Table 8. Within the metrics, for minADE metrics, smaller values indicate better more accurate perdictions, and for the rest metrics, the larger score means better performance. Within the results, we can see that forward prediction achieves much better performance compared to reverse prediction across all metrics, except similar performance in Angular speed, angular acceleration, distance to nearest object, and TTC. In these mentioned metrics. Note that the training time for forward prediction and reverse prediction are similar (10E6 and 15E6 steps respectively). The WOSAC results indicate that our BMT model is better at predicting future motion than predicting history motion. We will submit to the official WOSAC learderboard for further reference in the future.
@@ Appendix D Qualitative Results
Visualizations
In Fig. 6, we demonstrate six pairs of qualitative results by Adv-BMT. In different scenarios, our opponent behavior exihibits in different category of safety-critical driving patterns. This means that our augmented opponent agents have plausible interactions with realistic driving trajectories. Besides, in Fig. 7, we demonstrate diverse collisions from just one single scenario. This is the power of Adv-BMT compares to other baseline work, which generates the same or similar adversarial behavior with the same driving log.
Figure 7: Diverse adversarial behaviors generated on one driving log by Adv-BMT.
Demo Video
We submit a video within our supplementary materials. Here we put visualizations with case studies through animated visualizations of Adv-BMT scenarios, which simulates different types of vehicles, pedestrians and bicycle agents.

### v2

@@ A.1 BMT Architecture
Preprocessing.
@@
The model takes real-world scenarios in the ScenarioNet format li2023scenarionetopensourceplatformlargescale . It performs a series of preprocessing steps on both static map features and dynamic agent trajectories. We compute the global bounds to extract a consistent map center and heading. Each map feature is decomposed into a sequence of vectorized segments, where each vector is represented by start and end points in 3D coordinates. These vectors are augmented with relative directional positions, segment headings, and lengths. Semantic attributes are binary indicators that encode feature types. Map feature types (e.g., lanes, crosswalks, broken lines, yellow lines, stop signs, etc.) are encoded as semantic indicators. Features are then centered at the map center. Traffic-light states are also encoded and aligned with timesteps. Similarly, agent trajectories are centered for agent-feature encoding, and position, heading, velocity, shape, and type are extracted over time to form temporal sequences. We also reorder the ego agent and target agents to fixed indices. We extract a 16-dimensional state vector at each timestep. This preprocessing ensures that all trajectories are represented in a consistent spatial frame and expressed in an egocentric coordinate system.
BMT Motion Tokenization.
@@
We formulate motion prediction as an autoregressive next-token prediction problem, where each motion token corresponds to a discretized control input over a fixed time interval. The tokenization process maps continuous agent motion—longitudinal acceleration and yaw rate—into discrete 2D bins. During tokenization, we consider candidate tokens sampled from a fixed grid; for each candidate, we simulate the resulting motion over a short duration and select the best-matching action by minimizing the contour-alignment error between the predicted agent footprint and the ground-truth pose (position and heading). We adopt a simplified bicycle model to parameterize agent motion using longitudinal acceleration and yaw rate within predefined bounds: acceleration $\in[-10,10]\,\mathrm{m/s^{2}}$ and yaw rate $\in[-\pi/2,\pi/2]\,\mathrm{rad/s}$. Given predicted motion-token sequences, the trajectory is reconstructed by mapping tokens back to acceleration and yaw-rate changes.
@@
In both forward and reverse directions, the same tokenizer is used to decode motion tokens into continuous trajectories. Given an initial state $\tau_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, forward decoding simulates the next state $\tau_{t+1}$ by applying a tokenized control action $z_{t}=(a_{t},\delta_{t})\in\mathcal{A}$, where $\mathcal{A}$ is the discrete token space; repeating this autoregressively over the predicted tokens reconstructs the full trajectory $\boldsymbol{\tau}_{t:t+T}$. For reverse decoding, given a known future state $\tau_{t+1}$, the model evaluates candidate tokens $z_{t}\in\mathcal{A}$, simulates their inverse dynamics with $\Delta t\to-\Delta t$, and selects the token that best reconstructs the preceding state $\tau_{t}$. Operating in both directions is a key distinction of our approach: forward prediction enables open-loop rollouts of future behaviors, whereas reverse prediction traces from a desired outcome (e.g., a collision state) back to plausible initiating actions.
BMT Motion Decoder.
@@
The decoder follows a GPT-like structure composed of stacked cross-attention blocks. Each block integrates three structured attention modules:agent-to-agent (A2A), agent-to-temporal (A2T), and agent-to-scene (A2S). Relational information across modalities is encoded via relation embeddings: agent-to-agent relation embedding, agent-to-time relation embedding, and agent-to-scene relation embedding, each capturing context-specific spatial/temporal cues. For token construction, we use several embeddings, including the agent-type embedding, agent-shape embedding, agent-ID embedding, and motion-token embedding (which embeds discrete control tokens). A continuous-motion feature embedding encodes acceleration and yaw-rate attributes. Auxiliary embeddings include the special-token embedding (to indicate sequence boundaries) and the reverse-prediction indicator embedding (to distinguish forward from reverse prediction modes). For each attention edge, a relative-relation embedding is computed using a Fourier encoder tancik2020fourierfeaturesletnetworks and added to the key and/or value vectors. The input agent tokens are progressively updated across layers by aggregating contextual features from other agents, their temporal histories, and relevant map elements.
A.2 BMT Motion Decoding
Input.
@@ A.2 BMT Motion Decoding
Prediction Head and Output.
@@
where $\phi(\cdot)$ denotes the GELU activation and $K^{2}$ is the number of discrete motion tokens (from a $K\times K$ acceleration–yaw bin grid). The resulting logits are used to predict the next motion token at each time step. During inference, we generate motion tokens using nucleus (top-$p$) sampling.
@@
Our model makes a prediction in the interval of 0.5 seconds. To simulate the effect of a motion token over a fixed time step $\Delta t=0.5\,\mathrm{s}$, we adopt midpoint integration based on a simplified bicycle model. In forward predictions, given a current state $\mathbf{s}_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, the model computes the next speed and heading as $v_{t+1}=v_{t}+a\cdot\Delta t$ and $\theta_{t+1}=\theta_{t}+\omega\cdot\Delta t$. The average speed and heading are then defined as $\bar{v}=\frac{v_{t}+v_{t+1}}{2}$ and $\bar{\theta}=\left(\frac{\theta_{t}+\theta_{t+1}}{2}\right)$. In reverse prediction, the process is reverted. In the backward direction, the process is inverted. Given a future state $\mathbf{s}_{t+1}$, the model enumerates all possible token candidates and inverts the dynamics: $v_{t}=v_{t+1}-a\cdot\Delta t$ and $\theta_{t}=\theta_{t+1}-\omega\cdot\Delta t$. The average quantities $\bar{v}$ and $\bar{\theta}$ are computed similarly and used to derive the previous position:
@@
A.3 BMT Training
Training Loss.
@@ A.4 Training Details
Appendix B Additional Experiment Results
Training Environment.
@@ Appendix C Quantitative Results
C.1 Ablation Study
Table 9: Realism Results.
Method $_{\text{avg}} SFDE$$_{\text{min}}$ $_{\text{avg}} SADE$$_{\text{min}}$ $_{\text{agent,min}} Coll$$_{\text{agent,avg}}$
Reverse 3.92 2.69 5.53 5.11 0.03 0.05
Reverse + Adv-init 7.17 5.56 7.68 7.12 0.13 0.16
Reverse + Adv-init + Filter 6.93 5.30 7.58 7.02 0.12 0.15
Table 10: Diversity Results.
[…]
Reverse 8.35 3.13 0.17 0.71 0.02
Reverse + Adv-init 9.89 3.82 0.22 0.64 0.13
Reverse + Adv-init + Filter 10.27 3.99 0.22 0.73 0.09
Reverse predictions with adversarial initializations exhibit greater deviation from the ground-truth data and yield improved diversity. This behavior is expected, since adversarial initializations modify the terminal positions and headings of selected agents, which propagates backward into more varied histories. Importantly, our rule-based filtering does not reduce diversity; rather, it preserves multimodality while improving realism metrics by removing physically implausible trajectories.
C.2 Waymo Open Sim Agents Challenge
@@
We evaluate BMT results on 400 WOMD validation scenarios using the Waymo Open Sim Agents Challenge (WOSAC) 2025 montali2023waymoopensimagents . Evaluation results are summarized in Table 11. For the metrics, smaller values of minADE indicate more accurate predictions, whereas larger values for the remaining metrics indicate better performance. From the results, we observe that forward prediction achieves much better performance than reverse prediction across all metrics, except for similar performance in angular speed, angular acceleration, distance to the nearest object, and TTC. Note that the training times for forward prediction and reverse prediction are similar (10E6 and 15E6 steps, respectively). The WOSAC results indicate that our BMT model is better at predicting future motion than historical motion.
Upon reviewing the collisions detected in reverse predictions with real initializations, we found that most occurred in crowded parking-lot areas, where clusters of parked vehicles or pedestrians are close together. These were flagged as collisions by WOSAC’s evaluation API, even though they may not represent meaningful agent–agent interactions. This explains why metrics such as ADE and FDE remain comparable across methods despite differences in collision scores.
The performance gap is primarily due to our training strategy: the model was pre-trained for forward prediction (around 800k steps) and then fine-tuned for bidirectional prediction (around 1.5M steps). We have included these details in Section A.4 (Training Details).
@@ Appendix D Qualitative Results
Visualizations.
Figure 9 presents six pairs of qualitative results generated by Adv-BMT. Across different scenarios, the adversarial agents exhibit a diverse range of safety-critical driving behaviors, demonstrating their ability to interact plausibly with realistic traffic participants. The visualizations illustrate that Adv-BMT can generate multiple distinct collision outcomes from a single driving log. This highlights a key advantage of Adv-BMT over baseline methods, which tend to produce identical or highly similar adversarial behaviors for the same input scenario.
Demo Video.
We submit a video within our supplementary materials. Here we provide visualizations with case studies through animated simulations of Adv-BMT scenarios, which include different types of vehicles, pedestrians, and bicycle agents. More demos are available at https://metadriverse.github.io/adv-bmt/.
Figure 9: Diverse adversarial behaviors generated by Adv-BMT.